a) The following table shows the profile, by the math section of the SAT Reasoning Test,...

a)

The following table shows the profile, by the math section of the SAT Reasoning Test, of admitted students at UCLA for the Fall 2011 semester.

SAT Reasoning Test—Math Section
700–800 600–699 500–599 400–499 200–399 Total
Admitted 6,611 3,981 1,388 385 8 12,373
Not admitted 6,622 13,876 9,890 4,974 1,363 36,725
Total applicants 13,233 17,857 11,278 5,359 1,371 49,098

Determine the probability of the following event. (Round your answer to the nearest .01.) HINT [Example 6.]

An applicant neither had a Math SAT of 400 or above nor was admitted.

b)The following table shows the profile, by the math section of the SAT Reasoning Test, of admitted students at UCLA for the Fall 2011 semester.

SAT Reasoning Test—Math Section
700–800 600–699 500–599 400–499 200–399 Total
Admitted 6,611 3,981 1,388 385 8 12,373
Not admitted 6,622 13,876 9,890 4,974 1,363 36,725
Total applicants 13,233 17,857 11,278 5,359 1,371 49,098

Determine the probability of the following event. (Round your answer to the nearest .01.) HINT [Example 6.]

An applicant had a Math SAT below 500.

c)Suppose that in a year, 5% of all light vehicles (SUVs, pickups, passenger cars, and minivans) sold in the United States were pickups and 65% were passenger cars. Moreover, a randomly chosen vehicle sold that year was twice as likely to be an SUV as a minivan. Find the associated probability distribution.

Outcome SUVs pickups passenger cars minivans
Probability
Answer #1

a) An applicant neither had a Math SAT of 400 or above nor was admitted =(6622+13876+9890)/49098 =0.62

b) An applicant had a Math SAT below 500 =(13233+17857)/49098=0.63

c)

Outcome SUVs pickups passenger cars minivans
Probability 0.2 0.05 0.65 0.1